OpenStudy (anonymous):

Iodine-131 is an isotope that is used to study the functioning of the thyroid gland. This isotope has a half-life (time required for half the material to erode) of approximately 5 days. If a patient is given an injection that contains 5 micrograms of iodine-131, what will be the iodine level in the patient after 9 days? Use the function A(t) = 5 (1/2) t/5

OpenStudy (whpalmer4):

\[A(t) = 5(\frac{1}{2})^{t/5}\]Evaluate \(A(9) = 5(\frac{1}{2})^{9/5}\)